A) You invest $800 on May 25, 2018. On what day will you have earned exactly $31.96 of interest if your investment earns simple interest at r = 6%? 


Select one:

a. January 25, 2019

b. January 22, 2019

c. January 23, 2019

d. January 24, 2019

B)

$10,000 was deposited into an investment account earning interest at a nominal rate of 9% compounded monthly for eight years. How much interest was earned in the fifth year? 


Select one:

a. 1468.72

b. 1270.42

c. 1303.11

d. 1342.76

Please, explain the solutions
Answer #1

Solution A:

Option A is correct.

According to simple interest formula

I = P*R*T/12

31.96 = 800*0.06*T/12

T = 7.99 ~ 8 months

Therefore, 8 months from May 25,2018 is January 25, 2019.

Solution B:

Option D is correct.

Explanation:

Amount for year 4 = 10,000 (1 + 0.09/12)^4*12

Amount for year 4 = 10,000 (1.0075)^48

Amount for year 4 = $14314.05

Total interest till year 4 = $14,314.05 - $10,000 = $4,314.05

Amount for year 5 = 10,000 (1 + 0.09/12)^5*12

Amount for year 5 = 10,000 (1.0075)^60

Amount for year 5 = $15,656.81

Total interest till year 5 = $15,656.81 - $10,000 = $5,656.81

Interest for fifth year = $5,656.81 - $4,314.05

Interest for fifth year = $1,342.76